two men A and B are facing a flat and hard wall as shown. when man A makes a clap, man B hears two claps.if the speed of sound is 320m/s.what is the time interval between the two claps heard by man B?
1
Expert's answer
2018-11-19T16:02:10-0500

d_1=600 m

d_2=√(〖600〗^2+(400)^2 )+400=1121 m.

The time interval between the two claps:

t=(d_1-d_1)/v=(1121-600)/320=1.63 s.